OpenStudy (anonymous):

solve each system by elimination 6q - r + 2s = 8 2q + 3r - s = -9 4q + 2r + 5s = 1

OpenStudy (precal):

Multiply -3 to 2nd equation -3(2q+3r-s=-9) -6q-9r+3s=27 6q-r+2s=8 -10r+5s=35 multiply 2 to equation 1 and then add that to 3rd equation.